Fermi National Laboratory
 

Accelerator Update for January 8, 2007 - January 10, 2007

 

Monday January 8

The day shift began with Operations monitoring a Recycler stash (369.4E10), with the Antiproton Source (Pbar) stacking (57.0E10), with Linac, Booster, Main Injector (MI), and Recycler operating normally, with MiniBooNE off due to absorber work, with NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

At 8:23 AM, the MI Power Supply (MIPS) interlock loop tripped off with no indication as to why. The trip would not reset. An expert investigated and found that the LCW supply for MI-60 south had been turned off. He turned it on and the loop reset.

Operations put the Tevatron (TeV) into shot setup at 8:38 AM.

At 8:46 AM, the MIPS interlock loop at MI-30 tripped off for the first of two times in five minutes due to a volts-to-ground fault. It reset okay.

At 9 AM, the NuMI horn tripped off on a stripline ground fault. It reset okay.

Operations established store 5166 at 10:41 AM with an initial luminosity of 251.79E30. This store set a NEW RECORD luminosity.

Pbar experts began a stacktail study at 1:19 PM.

At 1:30 PM, TeV store 5166 aborted due to a power supply trip (T:SF). It was a clean abort, no quench. Power supply experts investigated T:SF.

Operations began preparing CDF and D0 for an access at 1:42 PM.

At 2:31 PM, the Pbar Debuncher dipole bus tripped off on a ground fault indication. Operations called an expert.

CDF came out of access at 2:48 PM.

TeV experts completed their repairs on T:SF.

D0 came out of access at 3:25 PM.

The Recycler completed its last of two stashes during the day shift at 3:34 PM.

MI experts began a barrier bucket study at 3:38 PM.

Operations put the TeV through a dry squeeze at 4:09 PM.

A Pbar expert reported at 5:14 PM that he had completed his work on the Pbar dipole bus, D:IB. In the process of bringing D:IB up, Operations cycled the Accumulator busses. Unfortunately, that aborted the antiprotons (7mA) in the Accumulator.

Operations loaded the TeV with protons for TeV studies at 5:42 PM. These studies ended at 7:46 PM.

Another TeV study began at 8:17 PM.

The Recycler completed its last of two stashes during the evening shift.

At 11:29 PM, problems with the Recycler beam cooling required Operations to call in an expert. Also, the NuMI beam permit tripped off due to problems with the beam position.

Tuesday January 9

The midnight shift began with TeV experts conducting studies, with Operations monitoring the stash (159.2E10), with Pbar stacking (72.9E10), and with NuMI taking beam.

The TeV studies ended at 12:15 AM.

Operations put the TeV through a dry squeeze at 12:33 AM.

The Recycler expert reported at 2:04 AM that he had resolved the cooling problem.

The Recycler completed a stash around 2:30 AM.

Operations put the TeV into shot setup at 3:41 AM.

Operations established store 5171 at 5:59 AM with an initial luminosity of 187.4E30.

At 7:20 AM, the NuMI horn tripped off on a ground fault indication. It reset.

A MI expert began a four-hour slipstacking study at 11:20 AM.

The Recycler completed its last of four stashes for the day shift at 1:43 PM.

At 4:51 PM, a Linac RF station (LRF4) tripped off. Operators and Linac experts investigated and then reset the station.

The NuMI beam permit tripped off at 6:12 PM for the first of three times in about three hours. It reset each time.

At 10:52 PM, the Tevatron Electron Lens (TEL) vacuum interlock trip caused TEL1 to trip off. CDF put their detector in stand-by. Operators investigated and found nothing wrong with the vacuum. The TEL was back on within five minutes. CDF resumed normal operations.

The Recycler stashed its last of four transfers for the evening shift around midnight.

Denton Morris took these pictures of Comet McNaught from St. Charles on Tuesday evening. The comet is visible to the naked eye near the horizon at sunset in the western sky.

Wednesday January 10

The midnight shift began with Operations monitoring store 5171 (38.25E30) and the stash (312.7E10), with Pbar stacking (14.2E10), and with NuMI taking beam.

At 1:03 AM, the NuMI beam permit tripped due to a beam position problem. It reset okay.

At 2:31 AM, Operations reported that Pbar suffered a shutter problem during a Recycler stash. Operators figured out how to complete the stash. Operations informed an expert.

The Recycler completed its only stash for the midnight shift at 3:31 AM.

Operations terminated store 5171 at 5:01 AM and then, a few minutes later, put the TeV into shot setup.

At 7:10 AM, Operations reported that the NuMI beam permit tripped due to an indication of current regulation problems on three power supplies. Everything reset okay.

A Pbar expert began his investigation of the shutter at 6:58 AM. Later he reported that he had replaced the shutter hardware.

Operations established store 5172 at 7:21 AM with an initial luminosity of 217.8E30. This luminosity seemed a little low for the amount of antiprotons. TeV experts will investigate.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow TeV, MI, Pbar, Linac, and Booster experts to conduct studies.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/11/2007    email Fermilab